A 7 kg body moving at 12 m/s accelerates to a velocity of 18 m/s in 4s time. What is the magnitude of the force that causes this increase in velocity?


1
Expert's answer
2021-03-16T11:37:47-0400
"a=\dfrac{v-v_0}{t}=\dfrac{18\ \dfrac{m}{s}-12\ \dfrac{m}{s}}{4\ s}=1.5\ \dfrac{m}{s^2},""F=ma=7\ kg\cdot1.5\ \dfrac{m}{s^2}=10.5\ N."